OpenStudy (anonymous):

A quality control inspector examined 200 light bulbs and found 15 of them to be defective. At this rate, how many defective bulbs will there be in a lot of 22000?

OpenStudy (anonymous):

1650 is what i got

OpenStudy (anonymous):

defective i mean

OpenStudy (anonymous):

So if they find 15 out of every 200, that means they'll find 30 out of 400, 45 out of 600, 60 out of 800, right? If you keep going up until 22000, you'll get your solution. How do you get to 22,000? by multiplying each time top and bottom 15/200. So to get the solution, you divide 22,000 by 200 and you get a number. The same number you need to multiply the 15 by. 22000/200=110 So now, just multiply 15*110 and you get 1650.

OpenStudy (anonymous):

\[\frac{15}{200}= \frac{x}{22000}\]\[x=1650 \]